BLACKJACK

Basic Strategy - 4 deck

Hard Totals:

HIT ≤ 8. STAND on ≥ 17. HIT 9, 10, 11 except, DOUBLE: 9 vs. dealer 3 − 6, 10 vs. dealer ≤ 9, 11 vs. dealer ≤ 10. STAND on 12 vs. dealer 4 − 6, otherwise HIT. STAND on 13 − 16 vs. dealer 2 − 6, otherwise HIT.

Soft Totals:

HIT soft 17 or lower. STAND on soft 19 or higher. HIT soft 18 vs. dealer 9, 10, A.

Splitting:

NEVER SPLIT 4s, 5s, 10s. ALWAYS SPLIT As, 8s. SPLIT 2s, 3s vs. dealer 4 − 7. SPLIT 6s vs. dealer 3 − 6. SPLIT 7s vs. dealer 7 or lower. SPLIT 9s vs. dealer 2 − 6, 8, 9.

Soft Doubling:

If this is allowed, DOUBLE A, 2 and A, 3 vs. dealer 5, 6. DOUBLE A, 4 and A, 5 vs. dealer 4, 5, 6 DOUBLE A, 6 and A, 7 vs. dealer 3, 4, 5, 6

Surrender:

If this is allowed, SURRENDER 9, 7 and 10, 6 vs. dealer 9, 10, A. SURRENDER 9, 6 and 10, 5 vs. dealer 10. If early surrender (before dealer looks at hole card) is allowed, SURRENDER all hard 5 − 7 and 12 − 17 vs. dealer A. SURRENDER all hard 14 − 16 vs. dealer 10.

\underline{Counting}

Count 3, 4, 5, 6 as +1 each; all 10s (including face cards) as −1. Divide this running count by number of decks to get true count. Positive counts are favorable to player (i.e., more 3s − 6s removed from deck and fewer 10s removed.) It is also favorable to player to have a larger than normal proportion of As left in the deck.